# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
where the implicit constant depends only on $\epsilon$, $\tilde{U}$, $\mathbf{G}$ and $\rho$. In the following, we will fix the constants $a_{w_{i}},b_{w_{i}},c_{w_{i}}$ and $A_{w_{i}}>0$.
Lemma 6.9.
Let $i\in\bar{I}$ such that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0.$ Then
$\max\{\alpha(a_{1}):\alpha\in\Phi(\mathbf{F}_{w_{i}})\}>0.$
Proof.
The proof is similar to those of Lemma 4.2 and Lemma 5.5. Suppose on the contrary that for all $\alpha\in\Phi(\mathbf{F}_{w_{i}})$, $\alpha(a_{1})=0$. Let $\Omega_{\mathbf{F}_{w_{i}}}$ and $\Omega_{\mathbf{H}_{w_{i}}}$ be small open neighborhoods of identity in $\mathbf{F}_{w_{i}}(\mathbb{R})$ and $\mathbf{H}_{w_{i}}(\mathbb{R})$ respectively, and
$\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}:=\Omega_{\mathbf{F}_{w_{i}}}\cdot\Omega_{\mathbf{H}_{w_{i}}}.$
Note that $w_{i}\mathbf{H}_{w_{i}}(\mathbb{R})w_{i}^{-1}$ stabilizes every element in $V_{\beta_{0}}$. Now for any $p=f\cdot h\in\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}$ with $f\in\Omega_{\mathbf{F}_{w_{i}}}$ and $h\in\Omega_{\mathbf{H}_{w_{i}}}$, we have
$a_{t}\cdot p\cdot w_{i}^{-1}=(a_{t}fa_{-t})(a_{t}ha_{-t})a_{t}w_{i}^{-1}=fw_{i}^{-1}\cdot(w_{i}a_{t}ha_{-t}w_{i}^{-1})\cdot w_{i}a_{t}w_{i}^{-1}.$
The element $w_{i}a_{t}w_{i}^{-1}$ is rational and
$\delta(\rho(w_{i}a_{t}w_{i}^{-1})\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})=e^{\beta_{0}(w_{i}a_{t}w_{i}^{-1})}.$
This implies that
$\delta(\rho(a_{t}\cdot p\cdot w_{i}^{-1})\cdot\mathbb{Z}^{d})\leq\kappa\cdot e^{-\beta_{0}(w_{i}a_{-1}w_{i}^{-1})t}$
for any $t>0$ where the constant $\kappa$ depends only on $\Omega_{\mathbf{F}_{w_{i}}}$ and $w_{i}$. Let $\Omega_{\mathbf{P}_{0}}$ be a small neighborhood of identity in $\mathbf{P}_{0}$. Then one can deduce from the inequality above that for any point $p\Gamma\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot w_{i}^{-1}\Gamma$, the orbit $a_{t}\cdot p\Gamma$ diverges in $\mathbf{G}(\mathbb{R})/\Gamma$. Set $\Gamma_{w_{i}}=\Gamma\cap w_{i}^{-1}\Gamma w_{i}$. Then $\Gamma_{w_{i}}$ is commensurable with $\Gamma$ and $w_{i}^{-1}\Gamma w_{i}$, and we can conclude that for any $p\Gamma_{w_{i}}\in\Omega_{\mathbf{P}_{0}}\cdot\Omega_{R_{u}(\overline{\mathbf{P}}_{0})}\cdot\Gamma_{w_{i}}$, the orbit $a_{t}\cdot p\Gamma_{w_{i}}$ diverges in $\mathbf{G}(\mathbb{R})^{0}/\Gamma_{w_{i}}$. Similiar to Lemma 4.2 and Lemma 5.5, this contradicts the ergodic property of the flow $\{a_{t}\}_{t\in\mathbb{R}}$ on the homogeneous subspace $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\Gamma_{w_{i}}\cap\tilde{\mathbf{G}}(\mathbb{R})^{0}$, where $\tilde{\mathbf{G}}$ is the product of some $\mathbb{Q}$-simple factors $\mathbf{G}_{i}$ $(1\leq i\leq s)$ of $\mathbf{G}$ such that $\{a_{t}\}_{t\in\mathbb{R}}$ projects nontrivially into $\mathbf{G}_{i}(\mathbb{R})$ $(1\leq i\leq s)$. This completes the proof of the lemma. ∎
@@
Now we compute an upper bound for the Hausdorff dimension of the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$. In the computation, we project the subset $S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by the natural projection map
@@
and compute the Hausdorff dimension of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$. Since $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}$ is a local isometry, by the countable stability of Hausdorff dimension, we then obtain the upper bounds for $\dim_{H}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}^{-1}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))))$ and $\dim_{H}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$.
Let $\tilde{\Omega}$ be a bounded fundamental domain of $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We choose a bounded subset $\Omega_{\mathbf{H}_{w_{i}}}$ in $\mathbf{H}_{w_{i}}(\mathbb{R})$ and a bounded subset $\Omega_{\mathbf{F}_{w_{i}}}$ in $\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w_{i}}}\cdot\Omega_{\mathbf{F}_{w_{i}}}.$
Now suppose that $q\in\mathbf{G}(\mathbb{R})$ is a rational element with $w_{i}^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Let $\tilde{q}\in\mathbf{G}(\mathbb{R})$ such that $w_{i}^{-1}\tilde{q}$ is a representative of $w_{i}^{-1}q$ in $\tilde{\Omega}$ with
$w_{i}^{-1}\tilde{q}=w_{i}^{-1}q\gamma$
for some $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then one can check that $\tilde{q}$ is still a rational element in $\mathbf{G}(\mathbb{R})$ and $d(\tilde{q})=d(q)$ since $\Gamma$ preserves the lattice $\mathbb{Z}^{d}$. So by Corollary 2.3, we still have
$w_{i}^{-1}\tilde{q}\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Since $w_{i}^{-1}\tilde{q}$ is an element in the bounded fundamental domain $\tilde{\Omega}$, it implies that
$w_{i}^{-1}\tilde{q}\in\Omega_{\mathbf{H}_{w_{i}}}\cdot(\mathbf{F}_{w_{i}}(\mathbb{Q})\cap\Omega_{\mathbf{F}_{w_{i}}}).$
@@
The set $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ consists of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in\left(a_{-t_{k}}E_{w_{i}}(C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t_{k}}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q_{k}))a_{t_{k}}\right)w_{i}^{-1}q_{k}(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$d(q_{k})\leq C_{0}(e^{-\tau t_{k}}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}$
and $w_{i}^{-1}q_{k}\in\tilde{\Omega}$.
@@
Let $\delta>0$ and $B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ be the small open ball of radius $\delta$ centered at identity in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $C_{\delta}>0$ depending only on $\delta$ such that for any $R>0$
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}(R)\subset E_{w_{i}}(C_{\delta}\cdot R)\cdot\mathbf{H}_{w_{i}}(\mathbb{R}).$
@@
Let $x\in B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ and $y\in E_{w_{i}}(R)$. Then $x\cdot y\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We write
@@
where $u\in\mathbf{F}_{w_{i}}(\mathbb{R})$ and $v\in\mathbf{H}_{w_{i}}(\mathbb{R})$. Then we have
$\|\Psi_{w_{i}}(u)\|=$ $\|\rho_{\beta_{0}}(w_{i}uw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\rho(w_{i}uvw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|$
$=$ $\|\rho(w_{i}xyw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq C_{\delta}\cdot R$
for some constant $C_{\delta}>0$ depending only on $\delta>0$. This completes the proof of the lemma. ∎
Note that by definition, for any rational element $g\in\mathbf{G}(\mathbb{R})$, every element $x$ in the $\Gamma$-coset $g\Gamma$ of $g$ is also rational, and $d(x)=d(g)$. We will use this fact in the following lemma.
@@
Let $i\in\bar{I}$ and $L>0$. Then the set of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$ is equal to
$\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}.$
@@
Let $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in R_{u}(\overline{\mathbf{P}}_{0})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0}))$ with $w_{i}\cdot g$ rational and $d(w_{i}\cdot g)=L$. We write
@@
for some $u\in\Omega_{\mathbf{H}_{w_{i}}}$, $v\in\Omega_{\mathbf{F}_{w_{i}}}$ and $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})$. Then we have
$w_{i}\cdot g=(w_{i}\cdot u\cdot w_{i}^{-1})\cdot w_{i}\cdot v\cdot\gamma$
and $w_{i}\cdot v$ is rational with
$d(w_{i}\cdot g)=d(w_{i}\cdot v)=L.$
By Corollary 2.3, we know that $v\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. This completes the proof of the lemma. ∎
Remark 6.13.
@@
$\mathcal{S}(L):=\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w_{i}}},y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}),d(w_{i}\cdot y)=L\}$
$\mathcal{F}(L):=\{y\in\Omega_{\mathbf{F}_{w_{i}}}\cap\mathbf{F}_{w_{i}}(\mathbb{Q}):d(w_{i}\cdot y)=L\}.$
Now we construct open covers of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{i}(\tau))$ with arbitrarily small diameters. Here we use the right-invariant metric $d_{R_{u}(\overline{\mathbf{P}}_{0})}$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which induces a metric $d_{R_{u}(\overline{\mathbf{P}}_{0})/\Gamma}$ on the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$. Note that by Corollary 6.8, for any $\epsilon>0$ and $R>0$, we have
$\operatorname{Vol}(E_{w_{i}}(R))\lesssim_{\epsilon}R^{a_{w_{i}}+\epsilon}$
where the implicit constant depends only on $\epsilon$ and $i\in\bar{I}$. Let $\nu_{0}$ be the $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
@@
By Lemma 6.9, we have $\nu_{0}(a_{1})>0$.
Fix $\delta>0$. Let $L>0$ and $t\in\mathbb{N}$ such that
$L\leq C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}.$
Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ such that $d(q)=L$ and $w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ The following subset
$\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
@@
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))\left(a_{-t}\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w_{i}}\left(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q)\right)\cdot a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)\cdot a_{t}\right)\cdot\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L).$
Since every point $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)$ satisfies the property that $w_{i}\cdot g$ is rational and $d(w_{i}\cdot g)=L$, it follows again from Lemma 6.12 that
$\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathcal{S}(L)=\mathcal{S}(L).$
@@
$\left(a_{-t}E_{w_{i}}(C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/d(q))a_{t}\right)\cdot w_{i}^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
@@
$a_{-t}E_{w_{i}}\left(C_{\delta}\cdot C_{0}(e^{-\tau t}\cdot e^{\beta_{0}(w_{i}a_{-t}w_{i}^{-1})})^{\dim V_{\beta_{0}}}/L\right)a_{t}\cdot\mathcal{S}(L).$
@@
$\mathcal{P}_{t,L}=\bigcup_{d(q)=L\textup{ and }w_{i}^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})}\mathcal{P}_{t,L}(q).$

### v2

@@ 6. An upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}$
where the implicit constant depends only on $\epsilon,\Omega_{\beta_{0}},\mathbf{G}$ and $\rho$. In the following, we will fix the constants $a_{w},b_{w},c_{w}$ in Corollary 6.8 and $A_{w}$ for the bounded open subset $\Omega_{\beta_{0}}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$.
@@
Now we compute an upper bound for the Hausdorff dimension of the subset $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$ $(\tau\leq\beta_{0}(wa_{-1}w^{-1}))$. In the computation, we project the subset $S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi)$ into the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by the natural projection map
@@
and compute the Hausdorff dimension of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$. Since $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}$ is a local isometry, by the countable stability of Hausdorff dimension, we then obtain upper bounds for $\dim_{H}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}^{-1}(\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))))$ and $\dim_{H}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$.
Let $\tilde{\Omega}$ be a bounded fundamental domain of $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We choose a bounded subset $\Omega_{\mathbf{H}_{w}}$ in $\mathbf{H}_{w}(\mathbb{R})$ and a bounded subset $\Omega_{\mathbf{F}_{w}}$ in $\mathbf{F}_{w}(\mathbb{R})$ such that
$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w}}\cdot\Omega_{\mathbf{F}_{w}}.$
Without loss of generality, we may assume that
$\tilde{\Omega}\subset\Omega_{\mathbf{H}_{w}}\cdot\Omega_{\mathbf{F}_{w}}\subset\Omega_{\beta_{0}}.$
Now suppose that $q\in\mathbf{G}(\mathbb{R})$ is a rational element with $w^{-1}\cdot q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Let $\tilde{q}\in\mathbf{G}(\mathbb{R})$ such that $w^{-1}\tilde{q}$ is a representative of $w^{-1}q$ in $\tilde{\Omega}$ with
$w^{-1}\tilde{q}=w^{-1}q\gamma$
for some $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then one can check that $\tilde{q}$ is still a rational element in $\mathbf{G}(\mathbb{R})$ and $0pt(\tilde{q})=0pt(q)$ since $\Gamma$ preserves the lattice $\mathbb{Z}^{d}$. So by Corollary 2.3, we have
$w^{-1}\tilde{q}\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{F}_{w}(\mathbb{Q}).$
Since $w^{-1}\tilde{q}$ is an element in the bounded fundamental domain $\tilde{\Omega}$, it implies that
$w^{-1}\tilde{q}\in\Omega_{\mathbf{H}_{w}}\cdot(\mathbf{F}_{w}(\mathbb{Q})\cap\Omega_{\mathbf{F}_{w}}).$
@@
The set $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\rho}(\psi)^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$ consists of points $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ for which there exist a divergent sequence $\{t_{k}\}\subset\mathbb{N}$ and a sequence of rational elements $\{q_{k}\}\subset\mathbf{G}(\mathbb{R})$ such that
$g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in\left(a_{-t_{k}}E_{w}(e^{\lambda_{w}(\epsilon)\cdot t_{k}}/0pt(q_{k}))a_{t_{k}}\right)\cdot w^{-1}q_{k}(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$0pt(q_{k})\leq e^{\lambda_{w}(\epsilon)\cdot t_{k}}$
and $w^{-1}q_{k}\in\tilde{\Omega}$.
@@
Let $\delta>0$ and $B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ be the small open ball of radius $\delta$ centered at identity in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then there exists a constant $C_{\delta}>0$ depending only on $\delta$ and $\rho$ such that for any $R>0$
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w}(R)\subset E_{w}(C_{\delta}\cdot R)\cdot\mathbf{H}_{w}(\mathbb{R}).$
@@
Let $x\in B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)$ and $y\in E_{w}(R)$. Then $x\cdot y\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and we write
@@
where $u\in\mathbf{F}_{w}(\mathbb{R})$ and $v\in\mathbf{H}_{w}(\mathbb{R})$. Then we have
$\|\Psi_{w}(u)\|=$ $\|\rho_{\beta_{0}}(wuw^{-1})\cdot e_{V_{\beta_{0}}}\|=\|\rho_{\beta_{0}}(wuvw^{-1})\cdot e_{V_{\beta_{0}}}\|$
$=$ $\|\rho_{\beta_{0}}(wxyw^{-1})\cdot e_{V_{\beta_{0}}}\|\leq C_{\delta}\cdot R$
for some constant $C_{\delta}>0$ depending only on $\delta>0$ and $\rho$. This completes the proof of the lemma. ∎
Note that by definition, for any rational element $g\in\mathbf{G}(\mathbb{R})$, every element $x$ in the $\Gamma$-coset $g\Gamma$ of $g$ is also rational, and $0pt(x)=0pt(g)$. We will use this fact in the following lemma.
@@
Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and $L>0$. Then the set of elements $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ with $w\cdot g$ rational and $0pt(w\cdot g)=L$ is equal to
$\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w}},y\in\Omega_{\mathbf{F}_{w}}\cap\mathbf{F}_{w}(\mathbb{Q}),0pt(w\cdot y)=L\}.$
@@
Let $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))\in R_{u}(\overline{\mathbf{P}}_{0})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0}))$ with $w\cdot g$ rational and $0pt(w\cdot g)=L$. We write
@@
for some $u\in\Omega_{\mathbf{H}_{w}}$, $v\in\Omega_{\mathbf{F}_{w}}$ and $\gamma\in\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})$. Then we have
$w\cdot g=(w\cdot u\cdot w^{-1})\cdot w\cdot v\cdot\gamma$
and $w\cdot v$ is rational with
$0pt(w\cdot g)=0pt(w\cdot v)=L.$
By Corollary 2.3, we know that $v\in\mathbf{F}_{w}(\mathbb{Q})$. This completes the proof of the lemma. ∎
Definition 6.12.
@@
$\mathcal{S}(L):=\{(x\cdot y)(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})):x\in\Omega_{\mathbf{H}_{w}},y\in\Omega_{\mathbf{F}_{w}}\cap\mathbf{F}_{w}(\mathbb{Q}),0pt(w\cdot y)=L\}$
$\mathcal{F}(L):=\{y\in\Omega_{\mathbf{F}_{w}}\cap\mathbf{F}_{w}(\mathbb{Q}):0pt(w\cdot y)=L\}.$
Now we construct open covers of $\pi_{R_{u}(\overline{\mathbf{P}}_{0})}(S_{\tau}^{c}\cap\mathcal{E}_{w,\epsilon}(\psi))$ with arbitrarily small diameters. Here we use the right-invariant metric $d_{R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})}$ on $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ which induces a metric on the quotient space $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})/(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$. Note that by Corollary 6.8, for any $\epsilon>0$ and $R>0$, we have
$\operatorname{Vol}(E_{w}(R))\lesssim_{\epsilon}R^{a_{w}+\epsilon}$
where the implicit constant depends only on $\epsilon$ and $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. Let $\nu_{0}$ be any $\mathbb{Q}$-root in $R_{u}(\overline{\mathbf{P}}_{0})$ such that
@@
Since the unstable horospherical subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\overline{\mathbf{P}}_{0})$, we have $\nu_{0}(a_{1})>0$.
Fix a small number $\delta>0$. Let $L>0$ and $t\in\mathbb{N}$ such that
$L\leq e^{\lambda_{w}(\epsilon)\cdot t}.$
Let $q$ be a rational element in $\mathbf{G}(\mathbb{R})$ such that $0pt(q)=L$ and $w^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}).$ The following subset
$\left(a_{-t}\cdot E_{w}\left(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q)\right)\cdot a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
@@
$B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta\cdot\exp(\nu_{0}(a_{-t})))\left(a_{-t}\cdot E_{w}\left(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q)\right)\cdot a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot B_{R_{u}(\overline{\mathbf{P}}_{0})}(\delta)\cdot E_{w}\left(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q)\right)\cdot a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
$\subset$ $\left(a_{-t}\cdot E_{w}\left(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L\right)\cdot a_{t}\right)\cdot\mathbf{H}_{w}(\mathbb{R})\cdot\mathcal{S}(L).$
Since every point $g(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$ in $\mathbf{H}_{w}(\mathbb{R})\cdot\mathcal{S}(L)$ satisfies the property that $w\cdot g$ is rational and $0pt(w\cdot g)=L$, it follows again from Lemma 6.11 that
$\mathbf{H}_{w}(\mathbb{R})\cdot\mathcal{S}(L)=\mathcal{S}(L).$
@@
$\left(a_{-t}E_{w}(e^{\lambda_{w}(\epsilon)\cdot t}/0pt(q))a_{t}\right)\cdot w^{-1}q\cdot(\Gamma\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R}))$
@@
$a_{-t}E_{w}\left(C_{\delta}\cdot e^{\lambda_{w}(\epsilon)\cdot t}/L\right)a_{t}\cdot\mathcal{S}(L).$
@@
$\mathcal{P}_{t,L}=\bigcup_{\begin{subarray}{c}0pt(q)=L\\ w^{-1}q\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\end{subarray}}\mathcal{P}_{t,L}(q).$